Having a hard time with the variance and standard deviation. Please help! Thank you!!!

Check my work mode : This shows what is correct or incorrect for the work you have completed so far. It does not indicate completion. Return to quer 3 Calculate the mean, the variance, and the standard deviation of the following discrete probability distribution. (Negative values should be indicated by a minus sign. Round intermediate calculetions to at least 4 decimal places. Round your final answers to 2 decimal places.) 4.16 points -44 -21 P(x x) 0.54 8.26 .13 .7 0 Answer is complete but not entirely correct. de 30.86 89 59 947 Mean Varlance deviation
Answer #1

mean=E(x)=sum(x*P(X=x))=-30.86

variance=E(x2)-E(x)*E(x)= 1176.46-(-30.86)*(-30.86)=224.1204

standard deviation=sqrt(variance)=sqrt(224.1204)=14.97

E(x2)=x2*P(X=x)=1176.46

following calcuation has been done for getting above answer

x P(X=x) x*P(X=x)   x2*P(X=x)
-44 0.54 -23.76 1045.44
-21 0.26 -5.46 114.66
-11 0.13 -1.43 15.73
-3 0.07 -0.21 0.63
sum= 1 -30.86 1176.46
variance= E(x2)-E(x)*E(x)= 224.1204
E(x2)=x2*P(X=x)